# Quantum Kernel Methods under Scrutiny: A Benchmarking Study

### v2

@@ D.5.2 Extended Correlation Analysis
D.5.3 Complementary Plots to Figure 12
In Sec. IV.5 we thoroughly investigate the underlying mechanisms that are eventually responsible for learning in PQK approaches. While comprehensive correlation analyses in Tabs. 1 and 2 provide profound insights, a deeper understanding of these findings follows from studying the distance measure as defined in Eq. (19). For Gaussian outer kernels we can additionally define the projected quantum circuit contributions $F_{\bm{\theta}}$ as given in Eq. (20). In the main text, we detail the corresponding discussion for the QFMNIST ($d=8$) dataset in case of the Gaussian outer kernel. The central illustration for this is provided in Fig. 12. In addition to these explanations we display the corresponding findings for the Matérn and RationalQuadratic outer kernels in Figs. A18 and A19, respectively.
Figure A18: Analogous illustration to Fig. 12 as discussed in Sec. IV.5 for the QFMNIST dataset with $d=8$ components and the Matérn outer kernel.
Figure A19: Analogous illustration to Fig. 12 as discussed in Sec. IV.5 for the QFMNIST dataset with $d=8$ components and the RationalQuadratic outer kernel.
@@ D.5.3 Complementary Plots to Figure 12
Appendix E Optimization of Trainable Parameters in Data Encoding Circuits
@@
Here, we provide a brief insight into the optimization of the variational parameters $\bm{\theta}$, which are generally incorporated for defining data encoding circuits as given in Eq. (3). As mentioned in Sec. III.1, these trainable parameters, if present, were randomly initialized with a fixed seed. To study the effectiveness of parameter optimization, we optimize the model corresponding to the best “ChebyshevPQC” configuration for both the two curves diff ($D=13$), cf. Fig. 8 (c) and the hidden manifold diff ($m=13$) datasets. For this purpose, the kernel-target-alignment (KTA) metric is maximized over the training set [21, 61]. We employ the respective sQUlearn implementation with the corresponding Adam optimizer.

### v3

@@ G.5.2 Extended Correlation Analysis
G.5.3 Complementary Plots to Figure 9
In Sec. 4.4 we thoroughly investigate the underlying mechanisms that are eventually responsible for learning in PQK approaches. While comprehensive correlation analyses in Tabs. 1 and 2 provide profound insights, a deeper understanding of these findings follows from studying the distance measure as defined in Eq. (15). For Gaussian outer kernels we can additionally define the projected quantum circuit contributions $F_{\boldsymbol{\theta}}$ as given in Eq. (16). In the main text, we detail the corresponding discussion for the QFMNIST ($d=8$) dataset in case of the Gaussian outer kernel. The central illustration for this is provided in Fig. 9. In addition to these explanations we display the corresponding findings for the Matérn and RationalQuadratic outer kernels in Figs. A21 and A22, respectively.
Figure A21: Analogous illustration to Fig. 9 as discussed in Sec. 4.4 for the QFMNIST dataset with $d=8$ components and the Matérn outer kernel.
Figure A22: Analogous illustration to Fig. 9 as discussed in Sec. 4.4 for the QFMNIST dataset with $d=8$ components and the RationalQuadratic outer kernel.
@@ G.5.3 Complementary Plots to Figure 9
Appendix H Optimization of Trainable Parameters in Data Encoding Circuits
@@
Here, we provide a brief insight into the optimization of the variational parameters $\boldsymbol{\theta}$, which are generally incorporated for defining data encoding circuits as given in Eq. (2). As mentioned in Sec. 3.1, these trainable parameters, if present, were randomly initialized with a fixed seed. To study the effectiveness of parameter optimization, we optimize the model corresponding to the best “ChebyshevPQC” configuration for both the two curves diff ($D=13$), cf. Fig. 8 (c) and the hidden manifold diff ($m=13$) datasets. For this purpose, the kernel-target-alignment (KTA) metric is maximized over the training set [23, 3]. We employ the respective sQUlearn implementation with the corresponding Adam optimizer.